using substituting method, solve 3x+3y=10 and 2x-y=2
Answer by addingup(3677)

3x+3y=10
2x-y=2 multiply all times 3 and add to equation above:
3x+3y=10
+
6x-3y=6
---------
9x = 16
x = 16/9 = 1 7/9